OpenStudy (bossimbacon):

Select the equation of the line in slope-intercept form: 4x 9y = -36 (Points : 4) Option A: 9y = 4x + 36 Option B: 9y = 4x 36 Option C: y = -4/9x - 4 Option D: y = 4/9x + 4 @alexandervonhumboldt2

OpenStudy (crystal_bliss):

Simplify 4x×9y to 36xy 36xy=−36 Divide both sides by 36 xy=− 36/36 Cancel 36 xy=−1 Divide both sides by y x=− 1/y

OpenStudy (denisaboichuk):

c